# Multimodal Medical Image Binding via Shared Text Embeddings

### v1

@@ 3.3. Adaptive Modality Balancing
(3) $p_{m}=\frac{\left(1/|D_{m}|\right)^{\beta}}{\sum_{m=1}^{M}\left(1/|D_{m}|\right)^{\beta}},$
where $\beta$ is the balancing hyperparameter that controls the degree of compensation for underrepresented modalities. Higher values of $\beta$ increase the sampling frequency of smaller datasets, while $\beta=0$ would result in uniform sampling regardless of dataset size. Second, we implement modality-specific learning rate scaling to optimize convergence across datasets of varying sizes. The learning rate $\eta_{m}$ for each modality $m$ is adjusted according to
(4) $\eta_{m}=\frac{\eta_{0}}{\sqrt{|D_{m}|}},$
@@
(5) $w_{m}=\frac{1}{\sqrt{|D_{m}|}}.$
These weights are applied to the corresponding modality-specific contrastive losses in the total objective function, ensuring that smaller datasets receive appropriate emphasis during training. The complete training objective integrates these balancing mechanisms into:
(6) $\mathcal{L}_{total}=\sum_{m=1}^{M}w_{m}\cdot\mathcal{L}_{CLIP}^{(m)}+\lambda\sum_{m_{1}\neq m_{2}}\mathcal{L}_{MSE}^{(m_{1},m_{2})},$
where the first term represents the weighted sum of modality-specific CLIP losses in Eq. (1), maintaining intra-modality alignment, and the second term aggregates MSE losses between all pairs of modalities $(m_{1},m_{2})$ to enforce alignment between the text embeddings. The hyperparameter $\lambda$ is the balancing coefficient. To preserve the well-established representational capabilities of the pre-trained CLIP-like models while efficiently adapting them to our multi-modal context, we leverage Low-Rank Adaptation (LoRA) (Hu et al., 2021) in optimizing Eq. (6).
3.4. Shared Embedding Space Knowledge Distillation
To consolidate the text embedding space into a single unified model, we propose Shared Embedding Space Knowledge Distillation (SESKD). The SESKD employs modality-specific text encoders as teachers, including ones for X-ray, Retina, Pathology, ECG, and CT, and distills their collective knowledge into a single student model based on BioClinicalBERT (Alsentzer et al., 2019). We choose BioClinicalBERT as it has been pre-trained specifically on biomedical and clinical text, providing domain-appropriate linguistic representations that align with the medical context of our imaging modalities. Our distillation process operates through two objectives. First, we apply a MSE loss to align the student’s text embeddings with each teacher encoder’s output, namely
(7) $\mathcal{L}_{\text{KD}}=\sum_{m=1}^{M}\text{MSE}(\mathbf{T}_{m},\mathbf{T}_{\text{student}}),$
@@ 3.4. Shared Embedding Space Knowledge Distillation
(8) $\mathcal{L}_{\text{Contrastive}}=-\sum_{i=1}^{N}\log\frac{\exp(\text{sim}(\mathbf{T}_{\text{student}},\mathbf{I}_{i})/\tau)}{\sum_{j=1}^{N}\exp(\text{sim}(\mathbf{T}_{\text{student}},\mathbf{I}_{j})/\tau)},$
@@
(9) $\mathcal{L}_{\text{SESKD}}=\alpha\cdot\mathcal{L}_{\text{KD}}+\mathcal{L}_{\text{Contrastive}},$
where $\alpha$ is a balancing hyperparameter that controls the relative importance of direct embedding alignment versus contrastive learning.
Table 1. Zero-Shot Performance Comparison Across Four Modalities (Retina, X-Ray, CT, Pathology) Using M3Bind and Baseline Models.
@@
MESIDOR(Decencière et al., 2014) ODIR200x31 CheXpert 200x5(Irvin et al., 2019) MIMIC-ECG(Gow et al., 2023) PTB-XL(Wagner et al., 2020)
@@
CLIP(Radford et al., 2021) 0.237 0.200 0.445 0.480 ConVIRT(Zhang et al., 2022) 21.30 19.03 17.65
BiomedCLIP(Zhang et al., 2023) 0.224 0.207 0.727 0.583 GLORIA(Huang et al., 2021) 23.20 15.79 44.01 ECG-CLIP 51.7 95.5 2.2 17.5
FLAIR-$\pi_{\text{naive}}$(Silva-Rodriguez et al., 2025) 0.545 0.442 0.447 0.470 PRIOR(Cheng et al., 2023) 34.15 29.83 34.38 MedBindNM(Gao et al., 2024) 50.2 93.9 1.9 18.2
FLAIR-$\pi_{\text{EK}}$(Silva-Rodriguez et al., 2025) 0.495 0.587 0.420 0.667 MGCA(Wang et al., 2022) 33.45 29.38 34.01 MedBindBD(Gao et al., 2024) 53.6 94.5 1.6 19.2
@@
SkinCancer(Kriegsmann et al., 2022) Camelyon (Veeling et al., 2018) SICAPv2 (Silva-Rodríguez et al., 2020) SkinTumor(Kriegsmann et al., 2022) CT-RATE(Hamamci et al., 2024) RAD-ChestCT(Draelos et al., 2021)
@@
CLIP(Radford et al., 2021) 0.05 0.62 0.39 0.10 CT-Net(Draelos et al., 2021) 0.603 0.631 0.581 0.239 0.544 0.587 0.540 0.285
BiomedCLIP(Zhang et al., 2023) 0.25 0.53 0.46 0.37 CT-CLIP(ZS)(Hamamci et al., 2024) 0.731 0.707 0.668 0.323 0.629 0.642 0.595 0.336
PLIP(Huang et al., 2023) 0.37 0.59 0.45 0.56 CT-CLIP(VF)(Hamamci et al., 2024) 0.756 0.738 0.705 0.353 0.650 0.659 0.659 0.349
QuiltNet(Ikezogwo et al., 2023) 0.45 0.65 0.39 0.58 CT-CLIP(CF)(Hamamci et al., 2024) 0.756 0.724 0.689 0.339 0.643 0.649 0.607 0.340
@@
4. Experiments
4.1. Pretrained CLIP-like Models Selected
@@
In our experiments, we select the following pre-trained CLIP-like models to train our M3Bind, namely MGCA (Wang et al., 2022) (specific to the X-Ray image), QuiltNet (Ikezogwo et al., 2023) (specific to the pathology image), FLAIR (Silva-Rodriguez et al., 2025) (specific to the retina image), and CT-CLIP (Hamamci et al., 2024) (specific to CT volumes). Regarding the ECG modality, since currently no related CLIP-like work is available, for subsequent fair comparisons, we follow the definition of the MedBind model (Gao et al., 2024) and train an ECG-CLIP model ourselves, which is then used for M3Bind training.
4.2. Datasets
We collected datasets spanning five distinct medical imaging modalities, namely X-ray, CT, retina, ECG, and pathology for M3Bind training and evaluation. For modalities with established CLIP-like models (X-ray, CT, retina, and pathology), we utilized data sets from original articles and maintained consistent preprocessing to ensure fair comparisons. For the ECG modality, which is used purely for evaluation, we follow the procedures described in MedBind (Gao et al., 2024). Specifically, for the training purpose of M3Bind, we use the dataset proposed in FLAIR(Silva-Rodriguez et al., 2025), which combines multiple retina image datasets and enhances them with expert knowledge. For evaluation, we used MESIDOR (Decencière et al., 2014) and an ODIR [^1] https://www.kaggle.com/datasets/andrewmvd/ocular-disease-recognition-odir5k test split, selected for their comprehensive coverage of the disease and their diverse imaging characteristics. For the X-ray image modality, we used MIMIC-CXR (Johnson et al., 2019a, b) for pre-training, which contains over 377,110 chest radiographs paired with radiology reports. Evaluation was performed using RSNA Pneumonia Detection (Shih et al., 2019) and CheXpert (Irvin et al., 2019), representing focused and multilabel diagnostic tasks. Following GLoRIA (Huang et al., 2021), we also constructed CheXpert5x200, a balanced subset containing 200 positive samples per disease class, to assess zero-shot classification capabilities in controlled class distributions. For the volumetric CT image modality, we used CT-RATE introduced in CT-CLIP (Hamamci et al., 2024) for both training and internal validation. This dataset consists of CT volumes paired with the corresponding radiology reports. For external validation, we employed RAD-ChestCT (Draelos et al., 2021), which provides an independent test set to evaluate generalization capabilities in different acquisition protocols and patient populations. Our pathology data consist of Quilt-1M(Ikezogwo et al., 2023), which combines the Quilt and OpenPath (Huang et al., 2023) datasets, totaling over a million whole slide image patches with corresponding textual descriptions for pre-training. For zero-shot evaluation, we evaluated performance in SkinCancer (Kriegsmann et al., 2022), Camelyon (Veeling et al., 2018), SICAPv2 (Silva-Rodríguez et al., 2020), and SkinTumor (Kriegsmann et al., 2022). Few-shot learning capabilities were evaluated using Camelyon (Veeling et al., 2018) and SICAPv2 (Silva-Rodríguez et al., 2020), which represent different organs and pathological conditions. For electrocardiogram (ECG) data, we utilized the MIMIC-ECG (Gow et al., 2023) dataset that contains 10-second, 12-lead recordings sampled at 500Hz, which we downsampled to 100Hz using a low-pass filter to maintain computational efficiency while preserving diagnostic features. Following MedBind (Gao et al., 2024), we generated paired textual descriptions for each ECG recording. An additional evaluation was performed on PTB-XL(Wagner et al., 2020) and ICBEB(Liu et al., 2018) to evaluate the generalization to the classification of cardiac abnormalities. Finally, to evaluate the cross-modal capabilities, we created MIMIC-PAIR by integrating MIMIC-CXR (Johnson et al., 2019a, b) and MIMIC-ECG(Gow et al., 2023) through hospital admission identifiers (hadm_id). When this identifier was unavailable, we matched cases according to subject_id, ensuring that chest radiographs and ECG recordings were temporally aligned within a 24-hour window.
4.3. Experimental Settings
@@
For all experiments, we maintained consistent preprocessing protocols as specified in the original papers for each modality-specific model (FLAIR, MGCA, CT-CLIP, ECG-CLIP, and QuiltNet) (Silva-Rodriguez et al., 2025; Wang et al., 2022; Hamamci et al., 2024; Ikezogwo et al., 2023). For the pre-training phase, we employed the AdamW optimizer with a base learning rate of 2e-5, incorporating warm-up and cosine decay schedules to ensure stable convergence. We configured batch sizes according to the complexity of the model and hardware constraints, namely 72 for most, with the exception of CT-CLIP (Hamamci et al., 2024) which required a reduced batch size of 8 due to its larger memory footprint. Text-modality alignment was performed with a batch size of 64 in all domains, using our adaptive sampling strategy to balance representation in datasets of different sizes. Training was carried out for 15,000 iterations with the balance factor $\lambda$ set to 10, prioritizing inter-modality text alignment before connecting the individual medical imaging modalities through the unified text space. In the SESKD training, we utilized a pretrained BioClinicalBERT model (Alsentzer et al., 2019) as our student text encoder. The training process was divided into two stages: first, we applied the knowledge distillation loss $\mathcal{L}_{\text{KD}}$ for 1,200 iterations with $\alpha=50$ to establish baseline alignment, then integrated the complete SESKD objective ($\mathcal{L}_{\text{SESKD}}$) for an additional 1,200 iterations to preserve both embedding similarity and contrastive relationships. All other optimization parameters remained consistent with the previous. All experiments were conducted on seven RTX 4090 GPUs, while the SESKD phase requires an additional RTX 4090 GPU.
4.4. Evaluations in Zero-Shot Manner
@@
For the Retina modality: Experiments on retina images were performed following FLAIR’s settings, which involved prompt mappings and disease-specific descriptions to enhance domain knowledge. Validation was performed on MESIDOR and ODIR datasets to test domain shift and novel class generalization. Comparisons were made with FLAIR-$\pi_{\text{naive}}$, FLAIR-$\pi_{\text{EK}}$, CLIP, and BiomedCLIP (Zhang et al., 2023). Our M3Bind achieved ACA scores (Zhao et al., 2019) of 0.564 and 0.543 on MESIDOR ($\pi_{naive}$ and $\pi_{EK}$, respectively) and 0.480 and 0.672 on ODIR200x3, demonstrating superior performance in both metrics.
@@
For the X-Ray modality: MIMIC-CXR pre-trained models were used to predict CheXpert labels, with performance evaluated on the CheXpert 5x200 test set using accuracy, F1 score, and precision metrics. The baseline comparisons shown in Table 1 included ConVIRT (Zhang et al., 2022), GLoRIA, MGCA, and PRIOR(Cheng et al., 2023). Our model achieved the highest scores, with an accuracy of 35.27, an F1-score of 30.71, and a precision of 34.96.
@@
For the Pathology modality: In addition to CLIP, BiomedCLIP, and QuiltNet, we also compared our approach with PLIP (Huang et al., 2023), another vision-language pre-training model for pathology. Apart from the SkinCancer dataset, we demonstrate performance improvements over QuiltNet on our other datasets, with overall performance exceeding any baseline.
For the CT modality: Internal validation was carried out on CT-RATE and external validation on RAD-ChestCT. Baseline comparisons included CT-Net(Draelos et al., 2021) and three variants of CT-CLIP (ZS, VF, CF). Our model achieved F1 scores of 0.742 in CLIP-RATE and 0.662 on RAD-ChestCT, reflecting high performance in both datasets.
@@
Table 2. Few-Shot classification performance (Accuracy %) comparison of M3Bind across different modalities in terms of accuracy (mean $\pm$ standard deviation) against modality-specific state-of-the-art models on ten datasets spanning five distinct medical imaging modalities. Evaluations are conducted with limited training examples to assess cross-modal transfer capabilities. Red and blue highlighting indicates the best and second-best performance, respectively.
@@
CLIP(Radford et al., 2021) $0.299\pm 0.02$ $0.420\pm 0.03$ $0.425\pm 0.04$ $0.686\pm 0.02$ $0.810\pm 0.03$ $0.823\pm 0.04$
BiomedCLIP(Zhang et al., 2023) $0.312\pm 0.03$ $0.430\pm 0.04$ $0.434\pm 0.03$ $0.712\pm 0.02$ $0.847\pm 0.03$ $0.846\pm 0.04$
ImageNet Init(Deng et al., 2009) $0.293\pm 0.02$ $0.413\pm 0.03$ $0.407\pm 0.05$ $0.615\pm 0.02$ $0.805\pm 0.03$ $0.813\pm 0.04$
FLAIR-$\pi_{\text{EK}}$(Silva-Rodriguez et al., 2025) $0.483\pm 0.05$ $0.601\pm 0.04$ $0.605\pm 0.03$ $0.784\pm 0.03$ $0.877\pm 0.04$ $0.875\pm 0.05$
M3Bind $0.467\pm 0.04$ $0.613\pm 0.03$ $0.620\pm 0.05$ $0.802\pm 0.03$ $0.890\pm 0.04$ $0.892\pm 0.05$
@@
RSNA(Shih et al., 2019) CheXpert(Irvin et al., 2019)
@@
ConVIRT(Zhang et al., 2022) $84.17\pm 0.77$ $86.92\pm 0.13$ $88.74\pm 0.36$ $85.02\pm 0.28$ $87.58\pm 0.53$ $88.21\pm 0.46$
GLoRIA(Huang et al., 2021) $84.12\pm 0.47$ $86.83\pm 0.53$ $89.13\pm 0.12$ $83.61\pm 0.52$ $87.40\pm 0.39$ $88.34\pm 0.12$
PRIOR(Cheng et al., 2023) $85.74\pm 0.36$ $87.08\pm 0.19$ $89.22\pm 0.16$ $86.16\pm 0.64$ $88.31\pm 0.20$ $88.61\pm 0.29$
MGCA(Wang et al., 2022) $85.80\pm 0.68$ $87.66\pm 0.21$ $89.30\pm 0.16$ $85.63\pm 0.33$ $87.65\pm 0.33$ $88.30\pm 1.48$
M3Bind $84.48\pm 0.12$ $87.54\pm 0.24$ $89.35\pm 0.43$ $88.07\pm 0.52$ $88.64\pm 0.27$ $88.98\pm 0.39$
@@
Camelyon(Veeling et al., 2018) SICAPv2(Silva-Rodríguez et al., 2020)
@@
CLIP(Radford et al., 2021) $80.38\pm 0.16$ $82.67\pm 0.19$ $82.80\pm 0.01$ $52.45\pm 2.41$ $62.24\pm 0.65$ $65.75\pm 0.16$
PLIP(Huang et al., 2023) $87.26\pm 0.23$ $87.48\pm 0.08$ $87.34\pm 0.01$ $65.76\pm 2.65$ $69.23\pm 0.43$ $73.00\pm 0.14$
QuiltNet(Ikezogwo et al., 2023) $87.62\pm 0.35$ $87.55\pm 0.03$ $87.48\pm 0.01$ $69.92\pm 1.02$ $74.14\pm 0.38$ $0.3675.31_{\pm}$ 0.36
M3Bind $87.65\pm 0.41$ $87.88\pm 0.37$ $87.63\pm 0.06$ $68.91\pm 1.12$ $74.56\pm 0.68$ $75.06\pm 0.10$
@@
CT-RATE(Hamamci et al., 2024) RAD-ChestCT(Draelos et al., 2021)
@@
CT-Net(Draelos et al., 2021) $63.12\pm 0.53$ $67.56\pm 0.61$ $69.75\pm 0.47$ $58.79\pm 0.73$ $63.86\pm 0.55$ $67.65\pm 0.63$
CT-CLIP(ZS)(Hamamci et al., 2024) $75.85\pm 0.68$ $81.17\pm 0.79$ $82.42\pm 0.57$ $65.37\pm 0.81$ $71.57\pm 0.42$ $74.58\pm 0.59$
CT-CLIP(VF)(Hamamci et al., 2024) $78.34\pm 0.49$ $82.99\pm 0.51$ $83.66\pm 0.73$ $68.02\pm 0.67$ $72.45\pm 0.72$ $75.42\pm 0.53$
CT-CLIP(CF)(Hamamci et al., 2024) $78.08\pm 0.65$ $82.85\pm 0.58$ $83.58\pm 0.79$ $67.81\pm 0.51$ $72.28\pm 0.43$ $75.13\pm 0.77$
M3Bind $78.77\pm 0.52$ $83.20\pm 0.46$ $83.39\pm 0.61$ $68.08\pm 0.41$ $72.35\pm 0.69$ $75.82\pm 0.57$
4.5. Evaluations in Few-Shot Manner

### v2

@@ 3.3 Adaptive Modality Balancing
$p_{m}=\frac{\left(1/|D_{m}|\right)^{\beta}}{\sum_{m=1}^{M}\left(1/|D_{m}|\right)^{\beta}},$ (3)
where $\beta$ is the balancing hyperparameter that controls the degree of compensation for underrepresented modalities. Higher values of $\beta$ increase the sampling frequency of smaller datasets, while $\beta=0$ would result in uniform sampling regardless of dataset size. We empirically set $\beta=0.5$ in our experiments. Second, we implement modality-specific learning rate scaling to optimize convergence across datasets of varying sizes. The learning rate $\eta_{m}$ for each modality $m$ is adjusted according to
$\eta_{m}=\frac{\eta_{0}}{\sqrt{|D_{m}|}},$ (4)
@@
$w_{m}=\frac{1}{\sqrt{|D_{m}|}}.$ (5)
These weights are applied to the total objective function, ensuring that smaller datasets receive appropriate emphasis during training. The complete training objective integrates these balancing mechanisms into:
$\mathcal{L}_{total}=\sum_{m=1}^{M}w_{m}\cdot\mathcal{L}_{CLIP}^{(m)}+\lambda\cdot w_{m_{1}}\cdot w_{m_{2}}\cdot\sum_{m_{1}\neq m_{2}}\mathcal{L}_{MSE}^{(m_{1},m_{2})}.$ (6)
where the first term represents the weighted sum of modality-specific CLIP losses in Eq. (1), maintaining intra-modality alignment, and the second term aggregates MSE losses between all pairs of modalities $(m_{1},m_{2})$ to enforce alignment between the text embeddings. The hyperparameter $\lambda$ is the balancing coefficient. To preserve the well-established representational capabilities of the pre-trained CLIP-like models while efficiently adapting them to our multi-modal context, we leverage Low-Rank Adaptation (LoRA) [15] in optimizing Eq. (6).
3.4 Shared Embedding Space Knowledge Distillation
To consolidate the text embedding space into a single unified model, we propose Shared Embedding Space Knowledge Distillation (SESKD). The SESKD employs modality-specific text encoders as teachers, including ones for X-ray, Retina, Pathology, ECG, and CT, and distills their collective knowledge into a single student model based on BioClinicalBERT [2]. We choose BioClinicalBERT as it has been pre-trained specifically on biomedical and clinical text, providing domain-appropriate linguistic representations that align with the medical context of our imaging modalities. Our distillation process operates through two objectives. First, we apply a MSE loss to align the student’s text embeddings with each teacher encoder’s output, namely
$\mathcal{L}_{\text{KD}}=\sum_{m=1}^{M}\text{MSE}(\mathbf{T}_{m},\mathbf{T}_{\text{student}}),$ (7)
@@ 3.4 Shared Embedding Space Knowledge Distillation
$\mathcal{L}_{\text{Contrastive}}=-\sum_{i=1}^{N}\log\frac{\exp(\text{sim}(\mathbf{T}_{\text{student}},\mathbf{I}_{i})/\tau)}{\sum_{j=1}^{N}\exp(\text{sim}(\mathbf{T}_{\text{student}},\mathbf{I}_{j})/\tau)},$ (8)
@@
$\mathcal{L}_{\text{SESKD}}=\mathcal{L}_{\text{KD}}+\mathcal{L}_{\text{Contrastive}},$ (9)
Table 1: Zero-Shot Performance Comparison Across Four Modalities (Retina, X-Ray, CT, Pathology) Using M3Bind and Baseline Models.
@@
MESIDOR[6] ODIR200x3LABEL:1 CheXpert 200x5[20] MIMIC-ECG[12] PTB-XL[45]
@@
CLIP[34] 0.237 0.200 0.445 0.480 ConVIRT[54] 21.30 19.03 17.65
BiomedCLIP[53] 0.224 0.207 0.727 0.583 GLORIA[16] 23.20 15.79 44.01 ECG-CLIP 51.7 95.5 2.2 17.5
FLAIR-$\pi_{\text{naive}}$[38] 0.545 0.442 0.447 0.470 PRIOR[3] 34.15 29.83 34.38 MedBindNM[9] 50.2 93.9 1.9 18.2
FLAIR-$\pi_{\text{EK}}$[38] 0.495 0.587 0.420 0.667 MGCA[46] 33.45 29.38 34.01 MedBindBD[9] 53.6 94.5 1.6 19.2
@@
SkinCancer[23] Camelyon [44] SICAPv2 [37] SkinTumor[23] CT-RATE[13] RAD-ChestCT[8]
@@
CLIP[34] 0.05 0.62 0.39 0.10 CT-Net[8] 0.603 0.631 0.581 0.239 0.544 0.587 0.540 0.285
BiomedCLIP[53] 0.25 0.53 0.46 0.37 CT-CLIP(ZS)[13] 0.731 0.707 0.668 0.323 0.629 0.642 0.595 0.336
PLIP[17] 0.37 0.59 0.45 0.56 CT-CLIP(VF)[13] 0.756 0.738 0.705 0.353 0.650 0.659 0.659 0.349
QuiltNet[19] 0.45 0.65 0.39 0.58 CT-CLIP(CF)[13] 0.756 0.724 0.689 0.339 0.643 0.649 0.607 0.340
@@
4 Experiments
4.1 Pretrained CLIP-like Models Selected
@@
In our experiments, we select the following pre-trained CLIP-like models to train our M3Bind, namely MGCA [46] (specific to the X-Ray image), QuiltNet [19] (specific to the pathology image), FLAIR [38] (specific to the retina image), and CT-CLIP [13] (specific to CT volumes). Regarding the ECG modality, since currently no related CLIP-like work is available, for subsequent fair comparisons, we follow the definition of the MedBind model [9] and train an ECG-CLIP model ourselves, which is then used for M3Bind training. We will primarily compare our approach with the selected CLIP-like models to demonstrate that our pre-training framework not only enables implicit alignment across different image modalities for these models, but also enhances their original performance.
4.2 Experimental Settings
@@
For all experiments, we maintained consistent preprocessing protocols as specified in the original papers for each modality-specific model (FLAIR, MGCA, CT-CLIP, ECG-CLIP, and QuiltNet) [38, 46, 13, 19]. For the pre-training phase, we employed the AdamW optimizer with a base learning rate of 2e-5, incorporating warm-up and cosine decay schedules to ensure stable convergence. We configured batch sizes according to the complexity of the model and hardware constraints, namely 72 for most, with the exception of CT-CLIP [13] which required a reduced batch size of 8 due to its larger memory footprint. Text-modality alignment was performed with a batch size of 64 in all domains, using our adaptive sampling strategy to balance representation in datasets of different sizes. Training was carried out for 15,000 iterations with the balance factor $\lambda$ set to 10, prioritizing inter-modality text alignment before connecting the individual medical imaging modalities through the unified text space. In the SESKD training, we utilized a pretrained BioClinicalBERT model [2] as our student text encoder. The training process was divided into two stages: first, we applied the knowledge distillation loss $\mathcal{L}_{\text{KD}}$ for 1,200 iterations to establish baseline alignment, then integrated the complete SESKD objective ($\mathcal{L}_{\text{SESKD}}$) for an additional 1,200 iterations to preserve both embedding similarity and contrastive relationships. All other optimization parameters remained consistent with the previous. All experiments were conducted on seven RTX 4090 GPUs, while the SESKD phase requires an additional RTX 4090 GPU.
4.3 Evaluations in Zero-Shot Manner
@@
For the Retina modality: Experiments on retina images were performed following FLAIR’s settings, which involved prompt mappings and disease-specific descriptions to enhance domain knowledge. Validation was performed on MESIDOR and ODIR datasets to test domain shift and novel class generalization. Comparisons were made with $\pi_{\text{naive}}, FLAIR$-$\pi_{\text{EK}}$, CLIP, and BiomedCLIP [53]. Our M3Bind achieved ACA scores [55] of 0.564 and 0.543 on MESIDOR ($\pi_{naive}$ and $\pi_{EK}$, respectively) and 0.480 and 0.672 on ODIR200x3, demonstrating superior performance in both metrics.
@@
For the X-Ray modality: MIMIC-CXR pre-trained models were used to predict CheXpert labels, with performance evaluated on the CheXpert 5x200 test set using accuracy, F1 score, and precision metrics. The baseline comparisons shown in Table 1 included ConVIRT [54], GLoRIA, MGCA, and PRIOR[3]. Our model achieved the highest scores, with an accuracy of 35.27, an F1-score of 30.71, and a precision of 34.96.
@@
For the Pathology modality: In addition to CLIP, BiomedCLIP, and QuiltNet, we also compared our approach with PLIP [17], another vision-language pre-training model for pathology. Apart from the SkinCancer dataset, we demonstrate performance improvements over QuiltNet on our other datasets, with overall performance exceeding any baseline.
For the CT modality: Internal validation was carried out on CT-RATE and external validation on RAD-ChestCT. Baseline comparisons included CT-Net[8] and three variants of CT-CLIP (ZS, VF, CF). Our model achieved F1 scores of 0.742 in CLIP-RATE and 0.662 on RAD-ChestCT, reflecting high performance in both datasets.
@@
Table 2: Few-Shot classification performance (Accuracy %) comparison of M3Bind across different modalities in terms of accuracy (mean $\pm$ standard deviation) against modality-specific state-of-the-art models on ten datasets spanning five distinct medical imaging modalities. Evaluations are conducted with limited training examples to assess cross-modal transfer capabilities. Red and blue highlighting indicates the best and second-best performance, respectively.
@@
CLIP $0.299\pm 0.02$ $0.420\pm 0.03$ $0.425\pm 0.04$ $0.686\pm 0.02$ $0.810\pm 0.03$ $0.823\pm 0.04$
[…]
ImageNet Init $0.293\pm 0.02$ $0.413\pm 0.03$ $0.407\pm 0.05$ $0.615\pm 0.02$ $0.805\pm 0.03$ $0.813\pm 0.04$
FLAIR- $\pi_{\text{EK}}$ $0.483\pm 0.05$ $0.601\pm 0.04$ $0.605\pm 0.03$ $0.784\pm 0.03$ $0.877\pm 0.04$ $0.875\pm 0.05$
OURS $0.467\pm 0.04$ $0.613\pm 0.03$ $0.642\pm 0.05$ $0.802\pm 0.03$ $0.890\pm 0.04$ $0.903\pm 0.05$
@@
RSNA CheXpert
@@
ConVIRT $84.17\pm 0.77$ $6.92\pm 0.13 88$$8.74\pm 0.36$ $85.02\pm 0.28$ $7.58\pm 0.53 88$$8.21\pm 0.46$
GLoRIA $84.12\pm 0.47$ $86.83\pm 0.53$ $89.13\pm 0.12$ $83.61\pm 0.52$ $7.40\pm 0.39 88$$8.34\pm 0.12$
PRIOR $85.74\pm 0.36$ $87.08\pm 0.19$ $89.22\pm 0.16$ $6.16\pm 0.64 88$$31\pm 0.20 88.$$8.61\pm 0.29$
MGCA $85.80\pm 0.68$ $87.66\pm 0.21$ $89.30\pm 0.16$ $85.63\pm 0.33$ $7.65\pm 0.33 88$$8.30\pm 1.48$
OURS $84.48\pm 0.12$ $87.54\pm 0.24$ $9.35\pm 0.43 88$$07\pm 0.52 88.$$64\pm 0.27 88.$$8.98\pm 0.39$
@@
Camelyon SICAPv2
@@
CLIP $80.73\pm 0.53$ $81.42\pm 0.21$ $82.07\pm 0.14$ $52.64\pm 1.91$ $62.79\pm 0.57$ $65.94\pm 0.43$
PLIP $87.26\pm 0.61$ $87.42\pm 0.23$ $87.45\pm 0.19$ $65.49\pm 0.95$ $68.97\pm 0.63$ $73.15\pm 0.71$
PLIP(FT) $87.50\pm 0.35$ $87.57\pm 0.18$ $87.53\pm 0.07$ $69.54\pm 0.47$ $73.18\pm 0.61$ $74.85\pm 0.53$
QuiltNet $87.51\pm 0.32$ $87.40\pm 0.13$ $87.27\pm 0.09$ $69.73\pm 0.78$ $74.19\pm 0.48$ $75.31\pm 0.36$
OURS $87.65\pm 0.41$ $87.88\pm 0.37$ $87.63\pm 0.06$ $68.91\pm 1.12$ $74.56\pm 0.68$ $75.54\pm 0.50$
@@
CT-RATE RAD-ChestCT
@@
CT-Net $63.12\pm 0.53$ $67.56\pm 0.61$ $69.75\pm 0.47$ $58.79\pm 0.73$ $63.86\pm 0.55$ $67.65\pm 0.63$
CT-CLIP(ZS) $75.85\pm 0.68$ $81.17\pm 0.79$ $82.42\pm 0.57$ $65.37\pm 0.81$ $71.57\pm 0.42$ $74.58\pm 0.59$
CT-CLIP(VF) $78.34\pm 0.49$ $82.99\pm 0.51$ $83.66\pm 0.73$ $68.02\pm 0.67$ $72.45\pm 0.72$ $75.42\pm 0.53$
CT-CLIP(CF) $78.08\pm 0.65$ $82.85\pm 0.58$ $83.58\pm 0.79$ $67.81\pm 0.51$ $72.28\pm 0.43$ $75.13\pm 0.77$
OURS $78.77\pm 0.52$ $83.20\pm 0.46$ $83.39\pm 0.61$ $68.08\pm 0.41$ $72.94\pm 0.69$ $75.82\pm 0.57$
4.4 Evaluations in Few-Shot Manner